### v2

# DAWN: Designing Distributed Agents in a Worldwide Network

Abstract
The rapid evolution of Large Language Models (LLMs) has transformed them from basic conversational tools into sophisticated entities capable of complex reasoning and decision-making. These advancements have led to the development of specialized LLM-based agents designed for diverse tasks such as coding and web browsing. As these agents become more capable, the need for a robust framework that facilitates global communication and collaboration among them towards advanced objectives has become increasingly critical. Distributed Agents in a Worldwide Network (DAWN) addresses this need by offering a versatile framework that integrates LLM-based agents with traditional software systems, enabling the creation of agentic applications suited for a wide range of use cases. DAWN enables distributed agents worldwide to register and be easily discovered through Gateway Agents. Collaborations among these agents are coordinated by a Principal Agent equipped with reasoning strategies. DAWN offers three operational modes: No-LLM Mode for deterministic tasks, Copilot for augmented decision-making, and LLM Agent for autonomous operations. Additionally, DAWN ensures the safety and security of agent collaborations globally through a dedicated safety, security, and compliance layer, protecting the network against attackers and adhering to stringent security and compliance standards. These features make DAWN a robust network for deploying agent-based applications across various industries.
Index Terms:
Large Language Model, Agent, Multi-agent system
I Introduction
Rapid advancements in large language models (LLMs)[1, 2, 3] have have sparked significant interest in LLM-based agents[4] which have the potential to free humans from repetitive tasks and significantly enhance productivity. New agents are being developed and their performance is improving on a regular basis. There are agents for many tasks such as coding[5, 6], web browsing[7], gaming, and many more.[8, 9]. Furthermore, multi-agent systems[10, 11, 12, 13, 14, 15, 16] have been proposed to leverage the collective intelligence and specialized profiles and skills of multiple agents. However, three major challenges arise that hinder their potential for broader, more impactful applications.
Figure 1: DAWN architecture showing its main components and capabilities.
The first challenge is how to discover and connect to agents distributed globally and accomplish complex tasks. Existing frameworks, such as LangGraph[14] and AutoGen[10] have bounded scope as they are typically designed to manage a limited number of agents. This constraint arises from the need to manually hard-code the names, URLs, and description of other agents into the prompts, which becomes increasingly difficult as the network of agents grows. More agents allow for a broader selection covering functional and non-functional capabilities. Additionally, integrating heterogeneous agents from different ecosystems requires developers to invest effort in developing wrappers and packages specific to each new framework as they evolve.
The second challenge is how to incorporate determinism into agentic applications, particularly when tasks require both flexible decision-making and predictable, reliable outcomes. While LLM-based agents excel in dynamic environments that exhibit flexibility, they falter in tasks requiring strict adherence to set execution graphs, predefined rules, and outcomes. Empirical evidence shows that manually designing a workflow rather than allowing agents to autonomously make decisions leads to more robust performance and cuts down on computational cost. Balancing the often orthogonal needs for autonomy and determinism is essential for the broader adoption of agent-based systems.
The third challenge is ensuring the safety and security of agentic applications. LLM agents, built using LLMs, are vulnerable to specific attacks such as prompt injection [17] and jailbreaking [18, 19]. Furthermore, in a global agentic framework, the use of dynamic capabilities can increase their exposure to widespread risks. For example, when LLM agents use third party APIs[20], malicious APIs can plan attacks and insert, substitute or delete key information. Depending on the agent’s application domain, such attacks could threaten physical security, financial security, or overall system integrity. This underscores the critical importance of applying robust safety and security measures in such frameworks.
To address these three issues, the Distributed Agents in a Worldwide Network (DAWN) framework is proposed. DAWN’s architecture includes key components such as the Principal Agent and Gateway Agents, which work together to enable agents to plan, compose, discover, connect, orchestrate, observe, safeguard, and communicate globally. Given a user request, the Principal Agent autonomously plans, produces, submits a list of tasks describing the derived capabilities to Gateway Agents,and then orchestrates these resources to execute the plan. (A Principal Agent is connected with a fixed of Gateway Agents in a fashion similar to sub-contractors associated with a principal contractor on a manufacturing contract.) Gateway Agents may be distributed globally and have the capability to register public and proprietary resources (tools, agents, and agentic applications). Upon request, the Gateway Agents search their registries for the most suitable resources to address the proffered task list. These selected resources are wrapped into callable sequences and returned to the Principal Agent. Fig. 1 illustrates the globally distributed nature of this collaboration through the Principal Agent and Gateway Agents.
DAWN offers multiple operational modes (and combinations of them) to orchestrate resources to accommodate tasks that require different level of determinism, as depicted in Fig. 2.:
Figure 2: DAWN offers levels of determinism and autonomy required for different modes of agentic operation
  • No-LLM Mode: This mode is best suited for tasks that require high predictability and reliability, leveraging traditional algorithms and tools to ensure consistent, repeatable results. Here, human operators manually design the workflow without using the Principal Agent and use only non-LLM tools retrieved from the Gateway Agents to maximize determinism.
  • Copilot Mode: In this mode, agents act as intelligent assistants, enhancing deterministic workflows with creative input. Human operators still design the workflow without using the Principal Agent. Tools, agents, and agentic applications retrieved from the Gateway Agents are employed in the intermediate steps.
  • LLM Agent Mode: In this mode, agents operate autonomously, making real-time decisions based on complex reasoning. Here, the Principal Agent defines the workflow and retrieves the necessary resources from the Gateway Agents autonomously. This mode is particularly effective in for complex tasks where business logic is fuzzy and there is tolerance for stochastic results.
The modes of operation are demonstrated in Section IV of the paper, where a human resources use case is used to exemplify the DAWN framework. DAWN implements safety, security, and compliance as a foundational layer within the architecture. It requires that all Gateway Agents meet minimum requirements across these three dimensions in order to participate. The DAWN framework has several key functions.
  • Multi-agent collaboration: DAWN leverages Gateway Agents to link agents globally. These Gateway Agents maintain registries that store agent profiles and intelligently match them to tasks. Thus, DAWN builds agentic applications by facilitating collaboration among agents, enabling them to discover, communicate, and coordinate across different systems and organizations. This set of Gateway Agents facilitates the execution of complex tasks by harnessing the individual expertise and collective intelligence of agents collaborating effectively.
  • Flexibility: DAWN introduces a versatile approach that adapts to a wide range of task requirements, from deterministic processes to dynamic decision-making. This adaptability allows for the seamless integration of traditional software tools, LLMs, and LLM agents, ensuring that the most appropriate method is used in each context.
  • Interoperability: DAWN is designed for seamless integration with a broad spectrum of tools, LLMs, and agents, but wrapping them as REST APIs in a unified format. It ensures interoperability across diverse ecosystems, which makes it adaptable to wider range of applications that require a blend of agentic processes and traditional software systems.
  • Connectivity: DAWN uses an open protocol to interact and communicate between distributed components. The Principal Agents maintains a channel that allows new Gateway Agents to request connections. These connections are accepted after due vetting. Similarly, communication between the Principal Agent and Gateway Agents occurs over encrypted channels using HTTPS. A similar arrangement exists between Gateway Agents and agentic resources.
  • Comprehensive Security and Compliance: The inclusion of a dedicated safety, security, and compliance Layer underscores the architecture’s commitment to maintaining high standards of data protection, user privacy, and regulatory compliance. This layer ensures that agentic operations within the framework adhere to stringent security protocols, protecting against unauthorized access and ensuring that the system meets industry-specific compliance requirements. Different organizations will demand varying levels of safety and security for handling agentic inputs and outputs and may specify these as a condition for participating in that organization’s implementation of the DAWN architecture.
II Related Works
II-A Large Language Models
Large Language Models have marked one of the most significant developments in artificial intelligence.Originally designed for tasks like machine translation and chat completion, state-of-the-art LLMs now demonstrate remarkable capabilities across a wide array of applications, including mathematics, coding, image analysis, image and video generation, speech recognition, and question answering. These successes have inspired researchers to propose the ”scaling law”[21], suggesting that continuous scaling of LLMs could eventually lead to the development of Artificial General Intelligence (AGI)[22]. Despite these impressive strides, one of the most pressing challenges is hallucinations—generating factually incorrect or nonsensical outputs[23, 24]. This issue arises because the internal knowledge within LLMs can be both limited and outdated, leading to errors that undermine their reliability and applicability in critical applications.
II-B LLM Agents
The concept of agents has been a staple in the field of artificial intelligence for decades[25]. However, it wasn’t until LLMs were used as reasoning engines within these agents that the concept truly began to flourish[26]. An LLM agent typically comprises three key components: a reasoning engine, action space, and memory unit. The reasoning engine, powered by an LLM, is equipped with advanced reasoning strategies such as Chain-of-Thought[27], ReAct[28], and Tree of Thought[29], enabling it to process and respond to complex queries. The action space is defined by a set of tools that the agent can utilize to interact with its environment and obtain feedback, facilitating dynamic decision-making. A memory unit plays a critical role by storing the reasoning trajectories, conversation history, and relevant context, allowing the agent to accumulate and apply knowledge over time. As a result, hundreds of such agents are being developed daily, with the expectation that they will soon transform traditional software and human processes, leading to enhanced productivity, improved user experiences, and greater operational efficiency.
II-C Multi-agent Systems
While individual agents have shown great promise, its true potential lies in the development of multi-agent systems. Human society has demonstrated that collaboration and teamwork often lead to superior outcomes, but the art of effective teamwork is complex and fraught with challenges[30]. Early efforts in building multi-agent systems have revealed tremendous potential[13, 11, 31, 32, 33, 12], but these works also highlight significant difficulties. Issues such as inefficient communication[12], increased computational costs, unreliable execution and communication sequences[34], and vulnerability to attacks[35] have underscored the need for a reliable platform to support agent collaboration. Additionally, scalability remains a critical concern: as the number of agents increases, introducing them to one another and ensuring that tasks are assigned to the most suitable agents becomes increasingly challenging. Addressing these concerns is essential for realizing the full potential of multi-agent systems in a wide range of applications.
Figure 3: DAWN reference architecture.
III DAWN Architecture Detail
The DAWN architecture is a modular framework designed to orchestrate collaboration of distributed agents. It supports agentic applications to be built with various operational modes, including No-LLM mode, Copilot mode, Agent mode, and combinations of them. It supports tasks ranging from deterministic workflows (e.g., job offers and background checks) to creative and autonomous processes (e.g., interview summaries and job description compliance remediation). Central to the Agent mode is the Principal Agent, which plans tasks and requests corresponding resources (tools, LLMs, and agents), and multiple Gateway Agents, each of which independently oversees resource retrieval, validation, and execution required for the tasks identified by the Principal Agent. Multiple agentic design patterns may be adopted to balance the performance and cost when operating in the Agent mode. DAWN uses a robust communication layer to facilitate seamless interaction between the Principal Agent, Gateway Agents, and registered agents distributed around the globe. A context layer maintains memory and task history, ensuring that the requisite information is exchanged efficiently via the communication layer. The safety, security, and compliance layer is responsible for the adherence of Principal and Gateway Agents to strict standards for the safe operation of agents during data and control interchange. Details of the architecture are shown in Fig. 3.
III-A Principal Agent
The Principal Agent acts as an autonomous planner-composer and central orchestrator within the DAWN framework. Its primary functions include creating a plan to execute customer requests, searching local resources (in the off chance some resources are available, a low probability occurrence), requesting additional resources from the Gateway Agents (the high probability occurrence), assessing resources returned from Gateway Agents for best fit, and delegating subtasks’ execution to the corresponding resources. Compared with the peer-to-peer communication pattern proposed by IoA[36] and AutoGen[10], using the Principal Agent as the central planner and orchestrator ensures an orderly and structured workflow that is efficient, reproducible, and easier to trace and restart if prematurely stopped. The Principal Agent first plans, seeks resources that meet its plan, and composes the task sequence using the retrieved resources. Orchestration, or the execution of the agentic resources, then begins with the planning using the cached information gathered during the compose phase to implement the sequence of resources required to fulfull the user’s intent. Other key components of the Principal Agent are as follows.
Reasoning, Planning, and Adaptive Decision-Making: The Principal Agent utilizes an LLM to understand the user’s intent, develop a plan, and decompose it into manageable subtasks. This feature is evident when DAWN operates in copilot or LLM mode. The Principal Agent’s planner uses ReAct [28], as the default reasoning strategy, which follows a cycle of reasoning, acting, and observing to actively move through subtasks. This dynamic approach enables the Principal Agent to adjust to complex and changing environments. For example, the user might ask the Principal Agent to plan a trip. The Principal Agent would first interpret the goal end-to-end and then create a plan with subtasks like booking a flight, reserving accommodation, and arranging local transport. If the flight turns out to be unavailable, the ReAct strategy allows the Principal Agent to search for alternative way such as train and bus, and modify the plan on the fly. Depending on the task requirements, users can select other reasoning strategies. For less dynamic tasks where computational cost is a priority, the Reasoning Without Observation (ReWOO) strategy [37] may be more suitable. ReWOO creates a comprehensive, end-to-end plan without adjusting it during execution. HuggingGPT[38] uses this strategy to process multi-modal data by calling machine learning models available on HuggingFace[39] and achieves good performance while minimizing LLM calls. Alternatively, the Tree-of-Thoughts (ToT) strategy [29] encourages the agent to explore multiple reasoning paths and self-evaluate them, offering more deliberate decision-making but at a higher computational cost. This approach is beneficial when performance is critical and cost is secondary. In Language Agent Tree Search (LATS)[40] and ToolLLM[41], the authors combine the ToT strategy with other techniques to achieve the best success rate that surpasses ReAct.
Local Resource Pool: The Principal Agent hosts a local pool for storing and caching resources retrieved from gateway Agent. Multiple common resources, such as a calculation tool and a web search tool, are permanently held in the local resource pool. References to resources retrieved from the Gateway Agents are also cached in the local resource pool using a Least Recent Used (LRU) strategy. This design Principal helps the Principal Agent to respond to simple or repetitive (sub)tasks without repeatedly reaching out to Gateway Agents and dramatically reduces the communication overhead. This resource pool plays a critical role in the orchestration phase of the task sequence.
Resource Requesting: After developing a plan, the Principal Agent summarizes the plan and its context—such as completed steps, conversation history, and user preferences—into queries and sends them to all connected Gateway Agents in an effort to retrieve the most appropriate resources to execute the plan in its entirety. If there are any suitable local resources, then the Principal Agent will not seek send out request for those.
Resource Execution: Gateway Agents are expected to apply their programmed discretion in searching, matching and retrieving the most applicable resources. When requests are sent to multiple Gateway Agents, it is only natural that duplicate resources or resources with overlapping capabilities are returned to the Principal Agent. The Principal Agent is responsible for analyzing the returned resource manifest files, assessing the quality of each resource, determining which agentic resource or resources may be applied to each task list step, and then retaining just those resources. When the Planning Agent begins execution of the task graph, it continues until it hits a node where a new search for resources is required to complete the remaining subtasks. (This situation can arise if during the compose phase of the process, the returned resources from the Gateway Agents only partially fulfill the customer need..) The Planning Agent then reaches out to the Gateway Agents again. This process repeats until the user’s request is fully addressed. It should be noted that Gateway Agents return references to agentic resources. For instance, some agents may actually execute on their preferred cloud service provider locations. In other cases, the agents may rely on the run-time environment of their host Gateway Agents. In both cases, only callable references to resources are passed between agents, Gateway Agents and Principal Agents. And so, when the Principal Agents executes resources from a Gateway Agent, it is in reality making an API call to the resource and passing the payload to the concerned Gateway Agent via JSON or similarly packaged property file format.
In many agentic implementations, the action space or toolset is limited to the local resource repository. This is typically due to three factors: 1) specialized agents naturally operate within a limited action space[42, 9, 43], 2) the context window of LLMs restricts the number of resources they can handle, and 3) if the agent has more than 20 tools to choose from, the accuracy in selecting the right tool drops as suggested by OpenAI[44]. (Our own experimentation found this number to be less than 10.) However, fully autonomous agents need to navigate real-world problems where the action space is vast, and numerous resources are available. DAWN’s rationale to use distributed Gateway Agents follows directly from this tenet. For example, if the Principal Agent needs to perform an action beyond an LLM’s capabilities—such as online search, booking a ticket, or writing code—it will request the relevant resource (e.g., a search tool, booking agent, or coding agent) from Gateway Agents. This flexibility enhances the Principal Agent’s autonomy and allows it to handle a broader range of use cases. However, this flexibility of this type comes with a cost. A Principal agent might get overlapping, conflicting, malicious or faulty resources. Dealing with these issues in software engineering will be quite a challenge. Overlapping resources require special handling to determine which resource is best in terms of functional as well as non-functional performance characteristics. In the case of malicious agents are introduced, whether inadvertent or otherwise, these agentic resources will render the entire distributed agentic application vulnerable to attack or worse, loss of faith. This is precisely why the safety and security of such applications are of crucial importance. Such faulty resources are an inherent part of the internet as well and people have been working to keep the defense line strong. This is the price we must pay for an open, participatory system. The alternate is a closed one and the trust that one provider will make it safe, secure, but inflexible.
III-B Gateway Agent
The Gateway Agent plays an essential role in connecting globally deployed resources (tools, agents, and agentic applications) with the Principal Agent. (There is a trivial case of human operators acting as the Principal Agent and using Gateway Agents for additional services under the No-LLM mode but the LLM-based applications are the central focus of this paper.) The Gateway Agent offers several services of which two are primary: a registry where resources are registered and an intelligent match-and-retrieve service for searching relevant resources given a query or task list. When the Principal Agent sends queries describing its need in natural language to the Gateway Agents, the Gateway Agents in turn search their registries and return the most appropriate resources to the Principal Agent. Each Gateway Agent may apply various filters and guardrails to select and trim the number of resources returned. Security measures such as access control and resource testing are integrated to reinforce safety of the system.
The Gateway Agent uses an open protocol, promoting decentralization and collaboration. Allowing different organizations and individuals to create and manage their own Gateway Agents leads to a global ecosystem of interconnected resources, encouraging innovation and diversity in problem-solving. It also adds a level of scalability and flexibility, as users would have access to a broader range of capabilities and knowledge through international cooperation. The key functions of Gateway Agents are listed below:
Resource Registration: The Gateway Agent hosts and maintains a registry where developers can register their resources (tools, agents, and agentic applications). A successfully registered resource will be accompanied with crucial details, such as the resource’s name, application programming interface (API), input and output schema, and API documentation. Information such as resource description, usage examples shall be included so that LLMs can understand what the resources are and how to use them through in-context learning. Gateway Agents may be equipped to run agentic resources within its confines or expect that agents have their own run-times.
When a resource is first registered with a Gateway Agent, the Gateway Agent has only limited information about it. It is screened for safety, security, compliance and other readily visible characteristics. Over time, the Gateway Agent also gathers utilitarian characteristics such as P50/P90 latency, cost, completion rate, etc. The registry may also include metrics such as success rate, average execution time, availability, and load capacity. These metrics allow the orchestrator to understand and optimize resource utilization and is then used to score, rate and select the best resource for a given user function.
Figure 4: Principal Agent and Gateway Agents workflow. For each user request, the Principal Agent first plans, produces a task list, and then requests resources from the Gateway Agents. The Gateway Agents return the most suitable resources to the Principal Agent, which is then used to first compose, and then orchestrate and execute the task list.
Resource Executor: For agents developed in different ecosystems to effectively communicate with the Principal Agent, execute tasks, and deliver results, they must adhere to a common protocol. In this case, the use of natural language specifications for agents and API descriptions can serve as the established common protocol. This protocol ensures seamless interaction and interoperability. As developers register their resources on the Gateway Agent, they are required to provide descriptions of their resources. These descriptions include details about the capabilities of the resources and instructions on how to utilize them.
Research has shown that Large Language Models (LLMs) can automatically communicate with RESTful APIs to execute tasks through them. RestGPT[45], for example, showcases this potential by leveraging the coding capabilities of LLMs to generate parsing code based on structured schemas. This approach automates both the generation of API calls and the parsing of responses and furthermore the LLM acts as a backup parser when necessary.
When incorporating agents from other frameworks and utilizing their manifest files, the Gateway Agent sends the manifest files to the Principal Agent. This enables the Principal Agent to communicate through a valid schema and ensures that the LLM can parse the response accurately based on the agent’s manifest file. By utilizing this mechanism, effective communication and parsing are maintained between the Principal Agent and other agents from different frameworks and allow for interoperability of such network of agents.
Resource Testing: Testing resources is crucial to ensure their proper functioning. Developers and providers who own the runtime environments are responsible for periodically testing and validating their resources. In addition to functionality testing, resources should be tested for connectivity, performance, security, and other factors [46]. For instance, it is important to evaluate how well a resource handles different levels of traffic. Various commercially available tools like Apache JMeter [47], Gatling [48], and Postman [49] provide the capability to test resources with a high volume of concurrent users.
If a resource provider fails to provide validation, the Gateway Agent will temporarily suspend the resource until the required tests are successfully completed, and the validation is provided. Likewise, resources managed by the Gateway Agents will undergo a policy check and alignment with guidelines through empirical testing and automated review [open ai store link]. If a resource fails to pass this validation process, indicating discrepancies between its description and actual behavior, the Gateway Agents will disqualify the agent from participating until a more accurate description is generated.
Resource Retrieval: A Gateway Agent handles queries from the Principal Agent, and searches its registry for the most relevant resources. Similar to traditional information retrieval methods [50], resource descriptions and examples may be treated like documents. Semantic search [51] may be used to identify the most relevant resources by converting both the queries and resource descriptions into vector embeddings and then performing similarity searches based on those embeddings. To further improve search accuracy, additional techniques like attribute filtering and keyword search are incorporated. Moreover, fine-tuning LLMs on query-resource pairs is another viable approach to enhance search results [52, 41]. While DAWN imposes some base functional and non-functional requirements on Gateway Agents along with stringent safety and security guardrails, its internals are left to the discretion of the Gateway owner. Some Gateway Agents may only have traditional software and classical data repositories, while others may be fully functioning LLM-driven Agents.
Multiple Gateway Agents: Organizations and developers around the globe will build and manage their own Gateway Agents following an open protocol. Each Gateway Agent offers unique resources proprietary to the entity in question (organization or developer). At any given time, the Principal Agent has a list of Gateway Agents it is connected with. The Principal Agent submits tasks and user intents to the listed Gateway Agents who then return the best resources available that may address all or a subset of the tasks or subtasks associated with the user request. The Principal Agent collects the search results and selects the highest-ranked (and most appropriate for the task) resources from them. This selection is based on factors like relevance, latency, cost, reputation, prior use, success and failure rates, and other metadata collected by the Gateway Agents about these agents over time. These metadata include information such as rating, latency, and performance of the registered agents. If the returned resources serve to complete the user tasks, then the Principal Agent’s task list is complete. If not, the cycle continues directed by the Principal Agent.
Gateway Agents may disconnect for maintenance or upgrades when idle or if its pending requests can be queued. Gateway Agents designed with memory and context mechanism to hold unaddressed user requests in a queue are able to connect and disconnect on their own schedules. A new Gateway Agent may join the Principal Agent’s list at any time but new Gateway Agents may only participate in a subsequent user request from the Principal Agent. This stands to reason as it might not have the context to address an ongoing task list. The Principal Agent continuously evaluates Gateway Agents based on the quality of the resources they provide. This evaluation helps assign weights and ratings to Gateway Agents and the resources they offer. Furthermore, users have the option to rate and assess the final execution of their request, including how each element was handled. These user ratings may also impact the ratings of Gateway Agents. The Principal Agent also uses its own memory module to learn from its own experiences. Fig. 4 illustrates the workflow between a Principal Agent and a Gateway Agents in full autonomous mode, incorporating most of the functionalities discussed above.
III-C Orchestration Layer
The Orchestration Layer in the Principal Agent, serves as a central component that enables the system’s versatility by supporting different operational modes, including No-LLM, Copilot, and hybrid modes that combine No-LLM, Copilot, and LLM Agent functionalities. In this hybrid mode, a human operator defines a workflow as an execution graph, where each node corresponds to a subtask, and the most suitable resource for each subtask is selected from those returned by the Gateway Agents. The Orchestration Layer is responsible for enforcing the rules, logic, and sequencing necessary to ensure that each subtask is executed in the correct order, adhering to the predefined workflow structure.
Additionally, the Orchestration Layer maintains the global state, tracking the progress of task execution across all nodes in the workflow and across all the participating Gateway Agents and global agentic resources. This global state ensures that each resource has the contextual information needed to perform its assigned subtask effectively. By maintaining and updating this global state, the system ensures that subtasks are executed with full awareness of prior actions and the broader context of the overall task.
III-D Communication Layer
The Communication Layer is responsible for managing the flow of information between various system components and resources, ensuring seamless interaction across the platform. It handles the messaging protocols that ensure messages are properly parsed, interpreted, and routed to their correct destinations. Two primary types of messages are exchanged among the system components and resources.
The first type is the query sent by the Principal Agent to the Gateway Agent. The payload initially contains the subtasks and their descriptions. This step also involves data sanitization and purposeful query generation to craft queries and tasks in a way that focuses on the specific information or task needed, without providing excessive user details. During task execution, if a resource fails, the Principal Agent will reach out to the Gateway Agents. The Principal Agent will provide the current subtask and its description as well as the completed subtasks and interaction history. This additional context enables the Gateway Agent to make informed decisions about which new resources to retrieve. In response, the Gateway Agent then returns a list of suitable resources along with manifest files that contain the full record of the functional and non-functional capabilities of the resources.
The second type of message is the execution command sent by the Principal Agent to the Gateway Agents’ resource executors. These message contain the necessary execution details, such as the resource endpoint and input parameters required for the task. A given resource executor may then respond with the result of the execution. In the event of an error during execution, the message will include an error notification to inform the Principal Agent, allowing for appropriate error handling and adjustments.
III-E Context Layer
The Context Layer is vital for enabling agentic operations by managing memory and task-related context. It provides three key components for memory management: a scratchpad to track the agent’s internal reasoning process, a message pool to record communication between the Principal Agent and the Gateway Agent, and a memory bank to store long-term data and user preferences, supporting long context, in-session, cross-session, multi-turn conversations resulting in personalized user experience. The Context Layer dynamically assembles prompts using data from both the message pool and memory bank, ensuring that the Principal Agent is aware of the overall task, task progress, and available resources.
Figure 5: Possible attacks on agentic networks and how DAWN mitigates such attacks. The Principal Agent and each of the Gateway Agents implement safety, security and compliance screening and filter to scrub agentic resources inputs and outputs.
III-F Security, Safety, and Compliance Layer
The Security, Safety, and Compliance Layer establishes a robust security foundation for the entire framework, addressing the safety, security, and compliance needs of the system. All frameworks and platforms that employ distributed resources must implement robust security systems to avoid risks that may be introduced when unsecured resources are connected to an enterprise’s gateway agent. While extensive research and practices have been published on traditional common security issues such as authentication, access control, privacy, and policy enforcement [53, 54, 55], an agentic network equipped with LLMs introduces additional security and privacy threats especially as resources worldwide begin collaborating. For LLM agents, two primary sources of threats exist.
First, there are vulnerabilities inherent to LLMs, such as hallucinations [56], attacks that leverage tuned instructional methods like jailbreaking [57, 58], and prompt injection [59]. Second, there are threats specific to agents collaborating in an agentic network like DAWN. A notable example is knowledge poisoning, where malicious data is integrated into the LLM’s knowledge base. PoisonedRAG [60] exemplifies this type of attack, where the corrupted knowledge base causes an agent to generate attacker-chosen responses to targeted questions. Another attack vector is functional manipulation. In the LLM agent’s workflow, after the Principal Agent takes an action, it processes the results and moves to the next step. An attacker can manipulate this process by inserting harmful prompts, thus directing the agent to perform unsafe tasks [61]. For instance, in a recruiting scenario the Principal Agent might retrieve a fraudulent review written by an attacker to recommend hiring an unsuitable candidate for a job. Agents are vulnerable to such unauthorized actions and similar manipulations.
There is extensive research dedicated to addressing both the inherent vulnerabilities of LLMs and attacks on agentic networks[62] . To protect DAWN framework against the knowledge poisoning, selecting reputable and transparent resources is essential. These resources must provide clear data usage policies and strong security records [63]. Tools and agents should also undergo third-party testing and sanitization, as demonstrated by ToolEmu [64], which proposes a framework for emulating tool execution through a language model. This approach allows scalable testing of the resources across various toolsets. Defense against functional manipulation begins with proactive security measures. For example, when employing third-party LLM agents, users should limit data sharing, particularly of sensitive personal information (PII). DAWN supports dynamic configurations where a co-pilot model integrates LLM-based and deterministic workflows, ensuring privacy and enterprise compliance often enforced by guardrails through algorithms designed to filter inputs and outputs[65, 66]. Incorporating deterministic steps into the workflow allows the enforcement of organizational policies and ensures user compliance through a LookUp tool, addressing potential instances where LLMs may bypass established guidelines. Additionally, increasing the reasoning capabilities of LLM agents is essential to counteract output manipulation. In the DAWN framework, the impact of output manipulation attacks is reduced by augmenting the Principal Agent with reasoning capabilities and access to a memory layer. Additionally, the modular architecture design enables the tracing of agentic attacks. Fig. 5 shows the attacks specific to agentic networks and DAWN’s approach in mitigating them.
The DAWN framework leverages Cisco’s expertise in safety and security to eliminate network attacks through traditional measures. Additionally, the outlined steps provide multiple layers of security within our agentic network, ensuring the system’s integrity, trustworthiness, and regulatory compliance.
Figure 6: Application of DAWN in an Human Resources recruiting use case in Agent + Copilot + No-LLM mode
IV Application of DAWN
The modular design of DAWN framework provides users with the flexibility to select the most appropriate working mode for their application, tailored to the specific needs of various use cases. As outlined in the introduction of this paper, the primary modes include No-LLM, Copilot, and LLM Agent mode. For scenarios requiring high determinism, users manually define plans, select suitable resources through the Gateway Agent, and assemble a No-LLM application within the orchestration layer, ensuring precise and predictable outcomes. In situations where uncertainty is present and decisions must be made dynamically based on real-time feedback, users manually construct a decision graph, incorporating an LLM to intelligently select which subgraph to execute based on evolving conditions. For cases with flexible workflow, users delegate the entire task to the Principal Agent, which autonomously plans, finds resources from the Gateway Agent, and constructs agentic applications. This adaptability ensures that the framework can efficiently handle a wide range of complex and diverse tasks.
To demonstrate DAWN’s capability, a walk-through example of a Human Resources (HR) recruiting use case is presented below. Finding the right candidate for a specific role in an organization involves multiple steps, most of which are repetitive and human-intensive, making them well-suited for LLM agents. When providing an HR application through DAWN, it is important to consider that the hiring process of each organization is optimized by HR experts according to that organization’s needs and policies. And it is crucial that not a single step be missed. Certain steps, such as offer negotiation with a candidate, require high security and are best handled by humans. Therefore, customers using DAWN for their agentic application are advised to use the framework in a hybrid mode.
Fig. 6 illustrates an HR example in a hybrid mode of Agent + Copilot + No-LLM. This hybrid mode includes a copilot that allows the customer to design the workflow manually, ensuring that none of the essential steps in the hiring process are neglected. The workflow is a directed graph designed by the customer, with most steps of the hiring process assigned to agentic nodes. These nodes leverage the full autonomous agent mode, where the Principal Agent requests resources from a Gateway Agent. In full autonomous mode, multiple specialized agents are seamlessly coordinated by the Principal Agent’s orchestration and the Gateway Agent’s resource retrieval. This setup promotes smooth cooperation and task efficiency. Additionally, the adaptive workflow allows the Principal Agent to dynamically adjust team members according to task progress. As it advances, the Principal Agent continuously communicates with the Gateway Agent for further assistance, ensuring the plan is regularly evaluated and updated if needed. This approach also supports iterative and traceable problem-solving, ensuring transparency and accountability throughout the entire process, making complex task management more effective.
Every step of the hiring process allocated to the agentic node involves the Principal Agent and Gateway Agent working together, showcasing the autonomous capabilities of DAWN’s framework. On the other hand, the No-LLM node demonstrates a more deterministic approach where a human takes over, primarily conducted through deterministic steps. This step includes offer negotiation and background checking, which involve candidate interaction and highly confidential and sensitive data that must be kept secure. By using the No-LLM mode, the DAWN framework allows for the integration of full determinism in the workflow when necessary. This HR use case was implemented with the listed resources to demonstrate the functionality of our framework, and leverages both expert knowledge and the autonomy of using LLM agents in applications requiring sophisticated task execution.
V Conclusion and Future Directions
The DAWN platform revolutionizes agentic applications by establishing a robust foundation for agent collaboration and communication. This framework offers a comprehensive strategy for integrating LLM-based agents into diverse applications. Its modular design, featuring key components like the Principal Agent and Gateway Agent, ensures that resources are optimally utilized, enabling seamless workflow management and global agent cooperation.
The DAWN platform’s contributions—hierarchical architecture engendering flexibility, scalability, interoperability, and stringent safety and compliance measures—position it as a transformative solution for small and large applications. By addressing the challenges of robustness and reliability in agent-based systems, DAWN paves the way for advanced utilization across various industries, from customer support to cybersecurity. The continued development and refinement of the DAWN framework is expected to push the boundaries of what is possible with LLM-based agents, fostering an interconnected ecosystem where collaboration and innovation thrive.
The innovative architecture of DAWN inspires future directions in agentic learning with abundant opportunities for agents to learn and adapt based on their interactions within the network. In DAWN, globally distributed agents have the potential to improve their effectiveness and presentation through manifest files, initially crafted by their developers to describe their resources. Gateway agents play a critical role here, gathering insights on agents’ functionality, latency, and security features, and adding this information as metadata to the agents’ manifest files. Through repeated interactions, the Principal agent learns from its experiences with different Gateway agents, optimizing its query patterns and enhancing its ability to select the most suitable Gateway agents for specific tasks. In turn, Gateway Agents learn about the agents that are registered with them. As each component gains experience and collaborates within the DAWN framework, they contribute to a collective improvement in efficiency, making task execution more cost-effective, secure, and time-efficient.
The evaluation of complex agentic workflows remains a dynamic field of research, with considerable focus on developing benchmarks and metrics for multi-agent applications. The lack of standardized benchmarks poses challenges, as does the assessment of individual network components in isolation—such as the Principal agent—due to the inherent complexity and non-determinism in multi-agent networks. Establishing clear metrics and success measures for multi-agent systems is essential, and DAWN’s architecture encourages further research and refinement in this field, potentially setting standards for agentic networks of the future.
VI Acknowledgment
This research is supported by Outshift at Cisco Systems. We extend our profound thanks to Reinaldo for his diligent review of our paper and for providing expertise that greatly contributed to enhancing the second version of the DAWN paper. We also express our sincere appreciation to Nihar Dandekar for sharing insights on potential solutions for agentic frameworks